Calculate (if possible) the values for the six trigonometric functions when tata= -450 degrees. HELP!!! THANKS!!!!
Answer by jsmallt9(3758)

Keys to this problem:
  • Understand that all angles in standard position that have the same terminal side will have the same sin, cos, etc. These are called co-terminal angles.
  • Co-terminal angles are multiples of 360 degrees apart from each other.
Combining these facts, we know that we can add 360 to or subtract 360 from any angle, multiple times if we want, and the sin, cos, etc. values will not change. So to find sin, cos, etc. of -450, we will add (or subtract) 360 as many times as it takes until we get an angle between 0 and 360. We want an angle between 0 and 360 because these are angles we are more familiar with.

Since -450 is negative, we need to be adding 360 to get an angle between 0 and 360:
-450 + 360 = -90
not yet...
-90 + 360 = 270
And we now have an angle between 0 and 360. We should recognize 270 as a special angle. And we should know what the sin, cos, etc. of 270 are:
sin(-450) = sin(270) = -1
cos(-450) = cos(270) = 0
tan(-450) = tan(270) = sin(270)/cos(270) = -1/0 which is undefined
csc(-450) = csc(270) = 1/sin(270) = 1/-1 = -1
sec(-450) = sec(270) = 1/cos(270) = 1/0 which is undefined
cot(-450) = cot(270) = cos(270)/sin(270) = 0/-1 = 0